# Proximal Interacting Particle Langevin Algorithms

### v1

@@ 4.5 Algorithm comparison
On the other hand, the bound for PIPGLA follows from first choosing $N$ so that the second term in Theorem 4.2 is $\mathcal{O}(\varepsilon)$, then $\lambda$ and $\gamma$ to counteract the dependence of $d_{x}$ on the fourth term while ensuring that the first term is at least $\mathcal{O}(\varepsilon)$. Finally, considering the values of $\lambda$ and $\gamma$, we select $n$ to ensure that the third term is $\mathcal{O}(\varepsilon)$.
@@
Evaluations of $\nabla g_{1}$ Evaluations of $\operatorname{prox}_{g_{2}}^{\lambda}$ Independent $1$d Gaussians
@@ 5.3 Bayesian neural network with sparse prior
(a) BNN with Laplace prior. Algorithm MYIPLA
(b) BNN with Normal prior. Algorithm IPLA
(c) BNN with Laplace prior. Algorithm IPLA
(d) BNN with Laplace prior. Algorithm PIPGLA
Figure 2: Histogram and density estimation of the weights for a randomly chosen particle from the final (500 steps) cloud of 100 particles.
@@ 5.3.1 Laplace vs Normal prior
Layer 1 Layer 2 Layer 1 Layer 2
@@ 5.4.1 Example 1
$U(\theta,\Phi,X)=\underbrace{-\log(C(\theta))+e^{-\theta}\sum_{l}\|X^{l}\|_{\text{tr}}}_{g_{2}(\beta,X)}+\underbrace{\sum_{l}\sum_{(i,j)\in\Omega^{l}}\frac{(Y_{ij}-X_{ij}^{l})^{2}}{2}}_{g_{1}(\beta,X)}.$
@@
$\frac{\|P_{\Omega^{\mathsf{c}}}Y_{\text{true}}-P_{\Omega^{\mathsf{c}}}Y_{K}^{N}\|_{F}^{2}}{\|P_{\Omega^{\mathsf{c}}}Y_{\text{true}}\|_{F}^{2}},$
where $P_{\Omega^{\mathsf{c}}}$ denotes the projection operator over the unobserved entries. The normalised error of the missing entries for the final particle cloud is $4.9\times 10^{-4}$, where we have considered $N=30$ particles and $K=500$ steps. The last value of the $\theta$ iterates is $\theta=0.1$, which provides a data-driven estimation of the regularisation parameter.
@@ A.1 Approximation of minimisers
Before proceeding with the proof of Theorem 4.1 and 4.2, we derive a result controlling the distance between the maximiser of $p_{\theta}(y)=\int_{\mathbb{R}^{d_{x}}}e^{-U(\theta,x)}\mathrm{d}x$, denoted by $\bar{\theta}_{\star}$, and the maximiser of $p_{\theta}^{\lambda}(y)=\int_{\mathbb{R}^{d_{x}}}e^{-U^{\lambda}(\theta,x)}\mathrm{d}x$, denoted by $\bar{\theta}_{\lambda,\star}$.
@@
$k_{\lambda}(\theta):=\int_{\mathbb{R}^{d_{x}}}e^{-U^{\lambda}(\theta,x)}\mathrm{d}x,\qquad\qquad k(\theta):=\int_{\mathbb{R}^{d_{x}}}e^{-U(\theta,x)}\mathrm{d}x$
@@
To obtain a bound of $\|\bar{\theta}_{\lambda,\star}-\bar{\theta}_{\star}\|$ in terms of $\lambda$, we first define the measure $\pi_{\lambda}^{1}\propto e^{-U^{\lambda}(\theta,x)}$ and observe that $\pi_{\lambda}^{1}$ is $\mu$-strongly log-concave, since $\pi_{\lambda}^{1}\propto e^{-U^{\lambda}(\theta,x)}$ and $U^{\lambda}$ is strongly convex by A4. Therefore, by a form of the Prékopa-Leindler inequality for strong convexity [Saumard and Wellner, 2014, Theorem 3.8], $\pi_{\lambda,\Theta}^{1}\propto e^{-K_{\lambda}(\theta)}=k_{\lambda}(\theta)$, is $\mu$-strongly log-concave, which results in
@@
$\|\nabla k_{\lambda}(\bar{\theta}_{\star})\|$ $=\Big{\|}\int_{\tilde{\Omega}}(\nabla_{\theta}g_{2}(\bar{\theta}_{\star},x)+\lambda h(\bar{\theta}_{\star},x)+\mathcal{O}(\lambda^{2}))e^{-U^{\lambda}(\bar{\theta}_{\star},x)}\mathrm{d}x\Big{\|}$
@@
$\|\nabla k_{\lambda}(\bar{\theta}_{\star})\|$ $\leq\Big{\|}\int_{\tilde{\Omega}}\nabla_{\theta}g_{2}(\bar{\theta}_{\star},x)\big{(}e^{-U^{\lambda}(\bar{\theta}_{\star},x)}-e^{-U(\bar{\theta}_{\star},x)}\big{)}\mathrm{d}x\Big{\|}+\big{(}\lambda\mathbb{E}_{X}[\|h(\bar{\theta}_{\star},X)\|]+\mathcal{O}(\lambda^{2})\big{)}k_{\lambda}(\bar{\theta}_{\star})$
@@
$\|\nabla k_{\lambda}(\bar{\theta}_{\star})\|$ $=\Big{\|}\int_{\tilde{\Omega}}\Big{(}\nabla_{\theta}g_{1}(\bar{\theta}_{\star},x)+\frac{\bar{\theta}_{\star}-\operatorname{prox}_{g_{2}}^{\lambda}(\bar{\theta}_{\star},x)}{\lambda}\Big{)}e^{-U^{\lambda}(\bar{\theta}_{\star},x)}\mathrm{d}x\Big{\|}$
@@
$\leq\int_{\tilde{\Omega}}\|\nabla_{\theta}U(\bar{\theta}_{\star},x)\|e^{-U^{\lambda}(\bar{\theta}_{\star},x)}\big{(}1-e^{-U(\bar{\theta}_{\star},x)+U^{\lambda}(\bar{\theta}_{\star},x)}\big{)}\mathrm{d}x+\big{(}\lambda\mathbb{E}_{X}[\|h(\bar{\theta}_{\star},X)\|]+\mathcal{O}(\lambda^{2})\big{)}k_{\lambda}(\bar{\theta}_{\star})$
@@ A.2 MYIPLA
$W_{2}(\pi_{\lambda,\Theta}^{N},\delta_{\bar{\theta}_{\star}})\leq W_{2}(\pi_{\lambda,\Theta}^{N},\delta_{\bar{\theta}_{\lambda,\star}})+\|\bar{\theta}_{\lambda,\star}-\bar{\theta}_{\star}\|\leq\sqrt{\frac{2d_{\theta}}{\mu N}}+\frac{\lambda}{\mu}\Big{(}\frac{\|g_{2}\|_{\textnormal{Lip}}^{2}}{2}A+B\Big{)}+\mathcal{O}(\lambda^{2}),$
@@
Since $\pi_{\lambda,\Theta}^{N}$ is smooth and satisfies the same assumptions required in Akyildiz et al. [2023, Proposition 3] to obtain a convergence bound for $W_{2}(\pi_{\lambda,\Theta}^{N},\delta_{\bar{\theta}_{\lambda,\star}})$, we have
$W_{2}(\pi_{\lambda,\Theta}^{N},\delta_{\bar{\theta}_{\lambda,\star}})\leq\sqrt{\frac{2d_{\theta}}{\mu N}}.$ (28)
@@
$\mathbb{E}[\|\theta_{n}^{N}-\bar{\theta}_{\star}\|^{2}]^{1/2}\leq$ $\sqrt{\frac{2d_{\theta}}{N\mu}}+\frac{\lambda}{\mu}\Big{(}\frac{\|g_{2}\|_{\textnormal{Lip}}^{2}}{2}A+B\Big{)}+e^{-\mu n\gamma}\bigg{(}\mathbb{E}[\|Z_{0}^{N}-z_{\star}\|^{2}]^{1/2}+\Big{(}\frac{d_{x}N+d_{\theta}}{N\mu}\Big{)}^{1/2}\bigg{)}$
@@
The first two terms are upper bounded by Proposition A.4, the second and third inequalities result from Proposition 4 and Proposition 5 of Akyildiz et al. [2023]. ∎
@@ A.3 PIPULA
$\mathbb{E}[\|\theta_{n}^{N}-\bar{\theta}_{\star}\|^{2}]^{1/2}\leq$ $\sqrt{\frac{2d_{\theta}}{N\mu}}+\frac{\gamma}{\mu}\Big{(}\frac{\|g_{2}\|_{\textnormal{Lip}}^{2}}{2}A+B\Big{)}+e^{-\mu n\gamma}\bigg{(}\mathbb{E}[\|Z_{0}^{N}-z_{\star}\|^{2}]^{1/2}+\Big{(}\frac{d_{x}N+d_{\theta}}{N\mu}\Big{)}^{1/2}\bigg{)}$
@@ A.4 PIPGLA
$W_{2}(\mathcal{L}({\theta}^{N}_{n}),\delta_{\bar{\theta}_{\star}})\leq\|\bar{\theta}_{\star}-\bar{\theta}_{\star,\lambda}\|+W_{2}(\delta_{\bar{\theta}_{\star,\lambda}},\pi_{\lambda,\Theta}^{N})+W_{2}(\pi_{\lambda,\Theta}^{N},\mathcal{L}({\theta_{n}^{N}})).$ (31)
We focus on the last term as we have already derived bounds for the first two terms. We begin by presenting some results that will be useful for proving a bound for $W_{2}(\pi_{\lambda,\Theta}^{N},\mathcal{L}({\theta_{n}^{N}}))$.
@@ A.4.1 Error bound for proximal gradient Langevin algorithm
Before deriving a bound for the last term in (31), we collect here a number of results adapted from Salim et al. [2019] which show convergence of the proximal gradient Langevin algorithm (PGLA) introduced in Salim et al. [2019] and recalled in Section 2.3.2. In particular, we derive convergence of the splitting scheme for general $\lambda$ (which includes as a special case $\lambda=\gamma$) when both $\nabla g_{1}$ and $\operatorname{prox}_{g_{2}}^{\lambda}$ can be computed exactly (which is a special case of the result in Salim et al. [2019] in the case $\lambda=\gamma$).
For convenience we consider the following decomposition of the PGLA update targeting a distribution $\pi_{\lambda}\propto\exp(-(g_{1}+g_{2}^{\lambda}))$ over $\mathbb{R}^{d}$
@@
We are going to derive a bound for $W_{2}(\mathcal{L}({X_{n+1}}),\pi_{\lambda})$, where $\mathcal{L}({X_{n+1}})$ denotes the distribution of the random variable $X_{n+1}$. For every $\pi$-integrable function $g:\mathbb{R}^{d}\to\mathbb{R}$, we define $\mathcal{E}_{g}(\pi)=\int g\mathrm{d}\pi$ and we denote $\mathcal{F}=\mathcal{E}_{g_{1}+g_{2}^{\lambda}}+\mathcal{H}$, where $\mathcal{H}$ is the negative entropy $\mathcal{H}(\pi)=\int\log\pi\;\mathrm{d}\pi$.
@@
Assume that $\mathbb{E}[\|\nabla g_{2}^{\lambda}(Y_{n})\|^{2}]\leq C$ for all $n\in\mathbb{N}$, where the expectation is w.r.t. the law of $\xi_{n}$.
@@
$2\gamma\big{[}\mathcal{E}_{g_{1}}(\mathcal{L}({Z_{n}}))-\mathcal{E}_{g_{1}}(\pi_{\lambda})\big{]}\leq(1-\gamma\mu)W_{2}^{2}(\mathcal{L}({X_{n}}),\pi_{\lambda})-W_{2}^{2}(\mathcal{L}({Z_{n}}),\pi_{\lambda}).$
@@
It is important to note, the above inequality is true for any $a$, $X_{n}$, and $Z_{n}$ where $Z_{n}=X_{n}-\gamma\nabla g_{1}(X_{n})$ (as deterministic vectors with appropriate dimension). Now, let $(a,X_{n})\sim\nu(\mathrm{d}a,\mathrm{d}x_{n})$ with marginal $\nu^{a}(\mathrm{d}a)=\pi_{\lambda,\star}(\mathrm{d}a)$. Taking conditional expectation w.r.t. $Z_{n}$ given $\sigma(a,X_{n})$, we obtain
@@
$\mathbb{E}\big{[}\|Z_{n}-a\|^{2}\big{]}\leq(1-\gamma\mu)\mathbb{E}_{\nu}\big{[}\|X_{n}-a\|^{2}\big{]}+2\gamma\big{(}\mathcal{E}_{g_{1}}(\pi_{\lambda})-\mathcal{E}_{g_{1}}(\mathcal{L}({Z_{n}}))\big{)}.$
@@
$W_{2}^{2}(\mathcal{L}({Z_{n}}),\pi_{\lambda})\leq(1-\gamma\mu)\mathbb{E}_{\nu}\big{[}\|X_{n}-a\|^{2}\big{]}+2\gamma\big{(}\mathcal{E}_{g_{1}}(\pi_{\lambda})-\mathcal{E}_{g_{1}}(\mathcal{L}({Z_{n}}))\big{)}.$
@@
$W_{2}^{2}(\mathcal{L}({Z_{n}}),\pi_{\lambda})\leq(1-\gamma\mu)W_{2}^{2}(\mathcal{L}({X_{n}}),\pi_{\lambda})+2\gamma\big{(}\mathcal{E}_{g_{1}}(\pi_{\lambda})-\mathcal{E}_{g_{1}}(\mathcal{L}({Z_{n}}))\big{)},$
@@
$\|y_{1}-a\|^{2}\leq\|y_{0}-a\|^{2}-2\lambda(g^{\lambda}(y_{0})-g^{\lambda}(a))+\lambda^{2}\|\nabla g^{\lambda}(y_{0})\|^{2}.$
@@
$\|y_{1}-a\|^{2}=\|y_{0}-a\|^{2}-2\lambda\langle\nabla g^{\lambda}(y_{0}),y_{0}-a\rangle+\lambda^{2}\|\nabla g^{\lambda}(y_{0})\|^{2},$
and using the convexity of $g^{\lambda}$ [Parikh and Boyd, 2014]
$\|y_{1}-a\|^{2}\leq\|y_{0}-a\|^{2}-2\lambda\big{(}g^{\lambda}(y_{0})-g^{\lambda}(a)\big{)}+\lambda^{2}\|\nabla g^{\lambda}(y_{0})\|^{2}.$
@@
$2\lambda\big{[}\mathcal{E}_{g_{2}^{\lambda}}(\mathcal{L}({Y_{n}}))-\mathcal{E}_{g_{2}^{\lambda}}(\pi_{\lambda})\big{]}\leq W_{2}^{2}(\mathcal{L}({Y_{n}}),\pi_{\lambda})-W_{2}^{2}(\mathcal{L}({X_{n+1}}),\pi_{\lambda})+\lambda^{2}C$
@@
$\|X_{n+1}-a\|^{2}$ $\leq\|Y_{n}-a\|^{2}-2\lambda\big{(}g_{2}^{\lambda}(Y_{n})-g_{2}^{\lambda}(a)\big{)}+\lambda^{2}\|\nabla g_{2}^{\lambda}(Y_{n})\|^{2}.$
Now, let $a$ be a random vector sampled from the distribution with density $\pi_{\lambda}$. Taking expectations in the previous expression and using the definition of the Wasserstein distance we obtain
$W_{2}^{2}(\mathcal{L}({X_{n+1}}),\pi_{\lambda})$ $\leq\mathbb{E}\big{[}\|Y_{n}-a\|^{2}\big{]}-2\lambda\big{(}\mathcal{E}_{g_{2}^{\lambda}}(\mathcal{L}({Y_{n}}))-\mathcal{E}_{g_{2}^{\lambda}}(\pi_{\lambda})\big{)}+\lambda^{2}\mathbb{E}[\|\nabla g_{2}^{\lambda}(Y_{n})\|^{2}]$
$\leq\mathbb{E}\big{[}\|Y_{n}-a\|^{2}\big{]}-2\lambda\big{(}\mathcal{E}_{g_{2}^{\lambda}}(\mathcal{L}({Y_{n}}))-\mathcal{E}_{g_{2}^{\lambda}}(\pi_{\lambda})\big{)}+\lambda^{2}C.$
Finally, taking the infimum over all couplings $Y_{n},a$ of $\mathcal{L}({Y_{n}}),\pi_{\lambda}$, it follows that
$W_{2}^{2}(\mathcal{L}({X_{n+1}}),\pi_{\lambda})\leq W_{2}^{2}(\mathcal{L}({Y_{n}}),\pi_{\lambda})-2\lambda\big{(}\mathcal{E}_{g_{2}^{\lambda}}(\mathcal{L}({Y_{n}}))-\mathcal{E}_{g_{2}^{\lambda}}(\pi_{\lambda})\big{)}+\lambda^{2}C.$

### v2

@@ 4.5 Algorithm comparison
On the other hand, the bound for PIPGLA follows from first choosing $N$ so that the first term in Theorem 4.2 is $\mathcal{O}(\varepsilon)$, then $\lambda$ and $\gamma$ to counteract the dependence of $d_{x}$ on the third term. Finally, considering the values of $\lambda$ and $\gamma$, we select $n$ to ensure that the second term is $\mathcal{O}(\varepsilon)$.
@@
Evaluations of $\nabla g_{1}$ Evaluations of $\operatorname{prox}_{g_{2}}^{\lambda}$ Indep. $1$d Gaussians
@@ 5.3 Bayesian neural network with sparse prior
(a) Laplace prior. Algorithm MYIPLA
(b) Normal prior. Algorithm IPLA
(c) Laplace prior. Algorithm IPLA
(d) Laplace prior. Algorithm PIPGLA
Figure 2: Histogram and density estimation of the weights of a BNN for a randomly chosen particle from the final (500 steps) cloud of 100 particles.
@@ 5.3.1 Laplace vs Normal prior
particles? Layer 1 Layer 2 Layer 1 Layer 2
@@ 5.4.1 Example 1
$U(\theta,\Phi,X)=\underbrace{-\log(C(\theta))+e^{-\theta}\sum_{l}\|X^{l}\|_{\text{tr}}}_{g_{2}(\beta,X)}+\underbrace{\sum_{l}\sum_{(i,j)\in\Omega^{l}}\frac{(Y_{ij}-\Phi X_{ij}^{l})^{2}}{2}}_{g_{1}(\beta,X)}.$
@@
$\frac{\|P_{\Omega^{\mathsf{c}}}Y_{\text{true}}-P_{\Omega^{\mathsf{c}}}Y_{K}^{N}\|^{2}}{\|P_{\Omega^{\mathsf{c}}}Y_{\text{true}}\|^{2}},$
where $P_{\Omega^{\mathsf{c}}}$ denotes the projection operator over the unobserved entries. The normalised error of the missing entries for the final particle cloud is $3.11\times 10^{-2}$, where we have considered $N=30$ particles and $K=500$ steps. The last value of the $\theta$ iterates is $\theta=0.1$, which provides a data-driven estimation of the regularisation parameter.
@@ A.1 Approximation of minimisers
Before proceeding with the proof of Theorem 4.1 and 4.2, we derive a result controlling the distance between the maximiser of $p_{\theta}(y)=\int_{{}^{d_{x}}}e^{-U(\theta,x)}\mathrm{d}x$, denoted by $\bar{\theta}_{\star}$, and the maximiser of $p_{\theta}^{\lambda}(y)=\int_{{}^{d_{x}}}e^{-U^{\lambda}(\theta,x)}\mathrm{d}x$, denoted by $\bar{\theta}_{\lambda,\star}$.
@@
$k_{\lambda}(\theta):=\int_{{}^{d_{x}}}e^{-U^{\lambda}(\theta,x)}\mathrm{d}x,\qquad\qquad k(\theta):=\int_{{}^{d_{x}}}e^{-U(\theta,x)}\mathrm{d}x$
@@
To obtain a bound of $\|\bar{\theta}_{\lambda,\star}-\bar{\theta}_{\star}\|$ in terms of $\lambda$, we first define the measure $\pi_{\lambda}^{1}\propto e^{-U^{\lambda}(\theta,x)}$ and observe that $\pi_{\lambda}^{1}$ is $\mu$-strongly log-concave, since $\pi_{\lambda}^{1}\propto e^{-U^{\lambda}(\theta,x)}$ and $U^{\lambda}$ is strongly convex by A4. Therefore, by a form of the Prékopa-Leindler inequality for strong convexity [Saumard and Wellner, 2014, Theorem 3.8], $\pi_{\lambda,\Theta}^{1}\propto e^{-K_{\lambda}(\theta)}=k_{\lambda}(\theta)$ is $\mu$-strongly log-concave, which results in
@@
[…]
@@
[…]
@@
[…]
@@
$\leq\int_{\tilde{\Omega}}\|\nabla_{\theta}U(\bar{\theta}_{\star},x)\|e^{-U^{\lambda}(\bar{\theta}_{\star},x)}\big{(}1-e^{-U(\bar{\theta}_{\star},x)+U^{\lambda}(\bar{\theta}_{\star},x)}\big{)}\mathrm{d}x$
@@ A.2 MYIPLA
$W_{2}(\pi_{\lambda,\Theta}^{N},\delta_{\bar{\theta}_{\star}})\leq W_{2}(\pi_{\lambda,\Theta}^{N},\delta_{\bar{\theta}_{\lambda,\star}})+\|\bar{\theta}_{\lambda,\star}-\bar{\theta}_{\star}\|\leq\sqrt{\frac{d_{\theta}}{\mu N}}+\frac{\lambda}{\mu}\Big{(}\frac{\|g_{2}\|_{\textnormal{Lip}}^{2}}{2}A+B\Big{)}+\mathcal{O}(\lambda^{2}),$
@@
Using a form of the Prékopa-Leindler inequality for strong convexity [Saumard and Wellner, 2014, Theorem 3.8], $\pi_{\lambda,\Theta}^{N}$ is $N\mu$-strongly log-concave. Since it is also smooth, we can apply Lemma A.8 of Altschuler and Chewi [2023] to obtain a convergence bound for $W_{2}(\pi_{\lambda,\Theta}^{N},\delta_{\bar{\theta}_{\lambda,\star}})$,
$W_{2}(\pi_{\lambda,\Theta}^{N},\delta_{\bar{\theta}_{\lambda,\star}})\leq\sqrt{\frac{d_{\theta}}{\mu N}}.$ (28)
@@
$\mathbb{E}[\|\theta_{n}^{N}-\bar{\theta}_{\star}\|^{2}]^{1/2}\leq$ $\sqrt{\frac{d_{\theta}}{N\mu}}+\frac{\lambda}{\mu}\Big{(}\frac{\|g_{2}\|_{\textnormal{Lip}}^{2}}{2}A+B\Big{)}+e^{-\mu n\gamma}\bigg{(}\mathbb{E}[\|Z_{0}^{N}-z_{\star}\|^{2}]^{1/2}+\Big{(}\frac{d_{x}N+d_{\theta}}{N\mu}\Big{)}^{1/2}\bigg{)}$
@@
The first two terms are upper bounded by Proposition A.4, the third and fourth inequalities result from Proposition 4 and Proposition 5 of Akyildiz et al. [2023]. ∎
@@ A.3 PIPULA
$\mathbb{E}[\|\theta_{n}^{N}-\bar{\theta}_{\star}\|^{2}]^{1/2}\leq$ $\sqrt{\frac{d_{\theta}}{N\mu}}+\frac{\gamma}{\mu}\Big{(}\frac{\|g_{2}\|_{\textnormal{Lip}}^{2}}{2}A+B\Big{)}$
$+e^{-\mu n\gamma}\bigg{(}\mathbb{E}[\|Z_{0}^{N}-z_{\star}\|^{2}]^{1/2}+\Big{(}\frac{d_{x}N+d_{\theta}}{N\mu}\Big{)}^{1/2}\bigg{)}$
@@ A.4 PIPGLA
$W_{2}(\mathcal{L}({\theta}^{N}_{n}),\delta_{\bar{\theta}_{\star}})\leq W_{2}(\delta_{\bar{\theta}_{\star}},\pi_{\Theta}^{N})+W_{2}(\pi_{\Theta}^{N},\mathcal{L}({\theta_{n}^{N}})).$ (31)
We begin by presenting some results that will be useful for proving a bound for $W_{2}(\pi_{\Theta}^{N},\mathcal{L}({\theta_{n}^{N}}))$.
@@ A.4.1 Error bound for proximal gradient Langevin algorithm
We collect here a number of results adapted from Salim et al. [2019] which show convergence of the proximal gradient Langevin algorithm (PGLA) introduced in Salim et al. [2019] and recalled in Section 2.3.2. In particular, we derive convergence of the splitting scheme for general $\lambda$ (which includes as a special case $\lambda=\gamma$) when both $\nabla g_{1}$ and $\operatorname{prox}_{g_{2}}^{\lambda}$ can be computed exactly (which is a special case of the result in Salim et al. [2019] in the case $\lambda=\gamma$).
For convenience we consider the following decomposition of the PGLA update targeting a distribution $\pi_{\lambda}\propto\exp(-(g_{1}+g_{2}^{\lambda}))$ over d
@@
We are going to derive a bound for $W_{2}(\mathcal{L}({X_{n+1}}),\pi_{\lambda})$, where $\mathcal{L}({X_{n+1}})$ denotes the distribution of the random variable $X_{n+1}$. For every $\pi$-integrable function $g:\mathbb{R}^{d}\to\mathbb{R}$, we define $\mathcal{E}_{g}(\pi)=\int g\mathrm{d}\pi$ and we denote $\mathcal{F}=\mathcal{E}_{g_{1}+g_{2}}+\mathcal{H}$, where $\mathcal{H}$ is the negative entropy $\mathcal{H}(\pi)=\int\log\pi\;\mathrm{d}\pi$. We also introduce the subdifferential of a convex function and its minimal section, since we will use them for our proofs.
Definition A.1 (Subdifferential and minimal section).
For any convex function $g:\mathbb{R}^{d}\to\mathbb{R}$, its subdifferential evaluated at $x$ is the set
$\partial g(x):=\{d\in\mathbb{R}^{d}\;|\;g(x)+\langle d,y-x\rangle\leq g(y)\forall y\in\mathbb{R}^{d}\}.$
Thanks to Bauschke and Combettes [2017, Proposition 16.4], we have that $\partial g(x)$ is a nonempty closed convex set. So the projection of 0 onto $\partial g(x)$, that is, the least norm element in the set $\partial g(x)$, is well defined, and we refer to this element as $\nabla^{0}g(x)$. Following Salim et al. [2019, Section 3.1], we name the function $\nabla^{0}g:\mathbb{R}^{d}\to\mathbb{R}$ the minimal section of $\partial g$.
@@
Assume that $\|\nabla^{0}g_{2}(x)\|^{2}\leq C$ for every $x\in\mathbb{R}^{d}$.
@@
$2\gamma\big{[}\mathcal{E}_{g_{1}}(\mathcal{L}({Z_{n}}))-\mathcal{E}_{g_{1}}(\pi)\big{]}\leq(1-\gamma\mu)W_{2}^{2}(\mathcal{L}({X_{n}}),\pi)-W_{2}^{2}(\mathcal{L}({Z_{n}}),\pi).$
@@
It is important to note, the above inequality is true for any $a$, $X_{n}$, and $Z_{n}$ where $Z_{n}=X_{n}-\gamma\nabla g_{1}(X_{n})$ (as deterministic vectors with appropriate dimension). Now, let $(a,X_{n})\sim\nu(\mathrm{d}a,\mathrm{d}x_{n})$ with marginal $\nu^{a}(\mathrm{d}a)=\pi(\mathrm{d}a)$. Taking conditional expectation w.r.t. $Z_{n}$ given $\sigma(a,X_{n})$, we obtain
@@
$\mathbb{E}\big{[}\|Z_{n}-a\|^{2}\big{]}\leq(1-\gamma\mu)\mathbb{E}_{\nu}\big{[}\|X_{n}-a\|^{2}\big{]}+2\gamma\big{(}\mathcal{E}_{g_{1}}(\pi)-\mathcal{E}_{g_{1}}(\mathcal{L}({Z_{n}}))\big{)}.$
@@
$W_{2}^{2}(\mathcal{L}({Z_{n}}),\pi)\leq(1-\gamma\mu)\mathbb{E}_{\nu}\big{[}\|X_{n}-a\|^{2}\big{]}+2\gamma\big{(}\mathcal{E}_{g_{1}}(\pi)-\mathcal{E}_{g_{1}}(\mathcal{L}({Z_{n}}))\big{)}.$
@@
$W_{2}^{2}(\mathcal{L}({Z_{n}}),\pi)\leq(1-\gamma\mu)W_{2}^{2}(\mathcal{L}({X_{n}}),\pi)+2\gamma\big{(}\mathcal{E}_{g_{1}}(\pi)-\mathcal{E}_{g_{1}}(\mathcal{L}({Z_{n}}))\big{)},$
@@
$\|y_{1}-a\|^{2}\leq\|y_{0}-a\|^{2}-2\lambda\left(g(y_{0})-g(a)\right)+\lambda^{2}\|\nabla g^{0}(y_{0})\|^{2}.$
@@
$\|y_{1}-a\|^{2}=\|y_{0}-a\|^{2}-2\lambda\langle\nabla g^{\lambda}(y_{0}),y_{0}-a\rangle+\lambda^{2}\|\nabla g^{\lambda}(y_{0})\|^{2}.$ (33)
Using that $y_{1}=y_{0}-\lambda\nabla g^{\lambda}(y_{0})$ we can write
$\langle\nabla g^{\lambda}(y_{0}),y_{0}-a\rangle=\langle\nabla g^{\lambda}(y_{0}),y_{1}-a\rangle+\lambda\|\nabla g^{\lambda}(y_{0})\|^{2}.$
Since $\nabla g^{\lambda}(y_{0})$ belongs to the subdifferential of $g(y_{1})$, i.e. $\nabla g^{\lambda}(x)\in\partial g(\text{prox}_{g}^{\lambda}(x))$ [Bauschke and Combettes, 2017, Proposition 16.44], we further have that
$\langle\nabla g^{\lambda}(y_{0}),y_{1}-a\rangle\geq g(y_{1})-g(a),$
from which we obtain
$-2\lambda\langle\nabla g^{\lambda}(y_{0}),y_{0}-a\rangle\leq-2\lambda\left(g(y_{1})-g(a)+\lambda\|\nabla g^{\lambda}(y_{0})\|^{2}\right).$
Recalling the definition of Moreau-Yosida approximation in Definition 2.2 we have that $g^{\lambda}(y_{0})=g(y_{1})+\|y_{0}-y_{1}\|^{2}/(2\lambda)$; plugging this into the equation above gives
$-2\lambda\langle\nabla g^{\lambda}(y_{0}),y_{0}-a\rangle$ $\leq-2\lambda\left(g^{\lambda}(y_{0})-g(a)\right)-2\lambda^{2}\|\nabla g^{\lambda}(y_{0})\|^{2}+\|y_{1}-y_{0}\|^{2}$ (34)
$=-2\lambda(g^{\lambda}(y_{0})-g(a))-\lambda^{2}\|\nabla g^{\lambda}(y_{0})\|^{2}.$
Finally, using Salim et al. [2019, Lemma 9] which states that $g^{\lambda}(x)\geq g(x)-\lambda\|\nabla^{0}g(x)\|/2$, where $\nabla^{0}g$ is the minimal section introduced in Definition A.1, and combining  (33) and (34) we have
$\|y_{1}-a\|^{2}$ $\leq\|y_{0}-a\|^{2}-2\lambda\left(g^{\lambda}(y_{0})-g(a)\right)\leq\|y_{0}-a\|^{2}-2\lambda\left(g(y_{0})-g(a)\right)+\lambda^{2}\|\nabla^{0}g(x)\|^{2}.$
@@
$2\lambda\big{[}\mathcal{E}_{g_{2}}(\mathcal{L}({Y_{n}}))-\mathcal{E}_{g_{2}}(\pi)\big{]}\leq W_{2}^{2}(\mathcal{L}({Y_{n}}),\pi)-W_{2}^{2}(\mathcal{L}({X_{n+1}}),\pi)+\lambda^{2}C$
@@
$\|X_{n+1}-a\|^{2}$ $\leq\|Y_{n}-a\|^{2}-2\lambda\big{(}g_{2}(Y_{n})-g_{2}(a)\big{)}+\lambda^{2}\|\nabla^{0}g_{2}(Y_{n})\|^{2}.$
Now, let $a$ be a random vector sampled from the distribution with density $\pi$. Taking expectations in the previous expression and using the definition of the Wasserstein distance we obtain
$W_{2}^{2}(\mathcal{L}({X_{n+1}}),\pi)$ $\leq\mathbb{E}\big{[}\|Y_{n}-a\|^{2}\big{]}-2\lambda\big{(}\mathcal{E}_{g_{2}}(\mathcal{L}({Y_{n}}))-\mathcal{E}_{g_{2}}(\pi)\big{)}+\lambda^{2}\mathbb{E}[\|\nabla^{0}g_{2}(Y_{n})\|^{2}]$
$\leq\mathbb{E}\big{[}\|Y_{n}-a\|^{2}\big{]}-2\lambda\big{(}\mathcal{E}_{g_{2}}(\mathcal{L}({Y_{n}}))-\mathcal{E}_{g_{2}}(\pi)\big{)}+\lambda^{2}C.$
Finally, taking the infimum over all couplings $Y_{n},a$ of $\mathcal{L}({Y_{n}}),\pi$, it follows that
$W_{2}^{2}(\mathcal{L}({X_{n+1}}),\pi)\leq W_{2}^{2}(\mathcal{L}({Y_{n}}),\pi)-2\lambda\big{(}\mathcal{E}_{g_{2}}(\mathcal{L}({Y_{n}}))-\mathcal{E}_{g_{2}}(\pi)\big{)}+\lambda^{2}C.$